OpenStudy (anonymous):

The temperature outside changed from 65°F to 41°F over a period of eight days. If the temperature changed by the same amount each day, what was the daily temperature change?

OpenStudy (anonymous):

65-41= 24 24/8 = 3 the daily temperature change was a -3 degrees each day

OpenStudy (anonymous):

it's not 3 though

OpenStudy (solomonzelman):

\(\huge {\color{purple} {\frac{65-41}{8}=?} }\)

OpenStudy (anonymous):

3. but it says it's wrong

OpenStudy (solomonzelman):

becuase it is \(\huge {\color{green} {-3} }\)

OpenStudy (anonymous):

oh O.O

OpenStudy (solomonzelman):

\(\huge {\color{green} {Because} }\) \(\huge {\color{green} {temperature} }\) \(\huge {\color{green} { increases} }\) \(\huge {\color{green} { 3°} }\) \(\huge {\color{green} {Each} }\) \(\huge {\color{green} {day.} }\)

OpenStudy (solomonzelman):

DECREASES

OpenStudy (anonymous):

Yes its -3 Makayla